OpenStudy (anonymous):

The product of complex numbers (8,9) and (5,-6) is ?

OpenStudy (jack1):

so u mean 8+9i and 5-6i, yeah? so product would be (8+9i)(5-6i) and i^2 = -1

OpenStudy (jack1):

if so = -54i^2 - 3i + 40 so = -54 x -1 - 3i + 40 = 94 - 3i